OpenStudy (anonymous):

One solution of the equation 6x^2-19x+3=0 is 1/6 What is the other solution?

OpenStudy (anonymous):

6x2 - 19x + 3 = 0 ( 6x - 1)( x - 3) = 0 x = 6/1 OR x = 3

OpenStudy (anonymous):

Omg thank you so much!!! :)

OpenStudy (anonymous):

your welcome